# Autonomous Navigation and Station-Keeping on Near-Rectilinear Halo Orbits

### v1

@@ 3.3.3 Measurement Update
$\hskip-2.0pt\boldsymbol{y}_{j}=\boldsymbol{E}\boldsymbol{\theta}_{j}+\boldsymbol{R}_{j},\;\;\boldsymbol{R}_{j}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{R}_{j}),\;\;\boldsymbol{E}=\begin{bmatrix}\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}\end{bmatrix}$. (29)
@@
$\hat{\boldsymbol{\theta}}_{j\mid j}$ $=\hat{\boldsymbol{\theta}}_{j\mid j-1}+\boldsymbol{K}_{j}\left(\boldsymbol{y}_{j}-\boldsymbol{E}\hat{\boldsymbol{\theta}}_{j\mid j-1}\right),$ (30b)
@@ 3.4 Measurement Collection for Horizon-Based Optical Navigation on Near-Rectilinear Halo Orbit
The choice of the onboard camera specifications directly impacts the number of limb points $m$ detected within an image, and therefore the quality of the position measurement itself. Furthermore, for a given position in space, if the field of view (FOV) is too narrow and the limb of the Moon does not fit within the frame of the image, the measurement may be unusable. The choice of camera FOV is particularly important on the NRHO due to the large variety in distance from the Moon at perilune and apolune. Figure 3 shows the range and the apparent diameter of the Moon as seen along the NRHO.
Figure 3: Range and apparent Moon diameter against osculating true anomaly along NRHO, shown for 10 superimposed revolutions
@@ 3.4.1 Effect of Field of View on Measurement
The least-squares solution of (7) is affected by the number of detected limb points $m$. One way to increase $m$ may be to use a camera with a larger number of pixels on its sensor; however, this requires larger onboard processing overhead. Another tuning knob is the camera’s FOV. A camera with an FOV that exactly matches the apparent diameter of the Moon will contain the largest possible $m$, assuming that the Moon is exactly centered along the boresight. However, in practice, the camera is pointed to the Moon using the state estimate $\hat{\boldsymbol{\theta}}$. Other operational considerations on attitude may further lead to scenarios where it is desirable to image the Moon without slewing the spacecraft too much.
@@ 3.4.2 Choice of Camera for Horizon-Based OPNAV on NRHO
When choosing the FOV, it is important to keep in mind that for the purpose of autonomous navigation, we seek to have reliable filter estimates near apolune, where station-keeping control maneuvers, when necessary, are executed. On one hand, there is an inherent degradation of observability as the range increases. The worsened observability is manifested in the slow variation of the Moon apparent diameter around apolune, as shown in Figure 3; as the image is taken further away from the Moon, a difference in range $\Delta r$ results in a rapidly diminishing difference in the apparent diameter. The degradation of the measurement quality is reported in previous works, including Figure 9 in [4]. Meanwhile, imaging at perilune, where we have the best observability, also has its downsides: first, from an operational perspective, this may lead to pointing requirements over short periods of time to ensure the Moon is within the frame of the camera. Second, a dead-reckoning prediction from perilune until the control location near apolune may be of insufficient accuracy. Third, even if a measurement collection opportunity may arise around apolune prior to control, a camera with a sufficiently large FOV to capture the full Moon near perilune will only lead to poor measurements with a small number of limb points $m$ near apolune. Thus, we choose an FOV sized to obtain measurements outside of regions along the orbit near perilune; to account for attitude uncertainties, we ensure the FOV is at least twice as large as the Moon’s apparent diameter from where the measurement is acquired.
@@ 4 Station-Keeping Control
There exists a multitude of controllers proposed for SK on LPOs [14]. In the context of L1 and L2 halo orbits, a popular SK controller is the so-called $x$-axis crossing control, which has been shown to perform well both in numerical simulations [16, 17] and in-flight [20]. A particularity about the $x$-axis crossing control scheme is that it has often been formulated as a multivariable root-finding problem, and rarely as a constrained minimization problem [33]. In this work, we introduce both the root-finding-based and the minimization problem-based formulations, point out their distinctions, and compare their numerical performances. Then, we propose a UT-based modification of the $x$-axis control that utilizes the filter’s covariance estimate together with the state estimate to improve the prediction accuracy of the targeted state, and thereby reduce the cumulative $\mathdollar\Delta V$.
Recent advancements in station-keeping on the NRHO also focused on tackling the phase-deviation issue associated with $x$-axis crossing control [19, 34]. To avoid the issue altogether, a model predictive control (MPC) scheme has also been proposed to provide full-state tracking under the same operational assumptions as $x$-axis crossing control [35], yielding similar cumulative $\Delta V$ and improved tracking performance. In this work, the conducted numerical experiments do not exceed $\sim 400$  $\text{\,}\mathrm{d}\mathrm{a}\mathrm{y}\mathrm{s}$, and thus are minimally impacted by phase deviation. The UT-based modification proposed here is applicable to any control scheme, including $x$-axis crossing control schemes with phase-deviation remedies [19, 34] or the MPC in [35].
In the rest of this Section, we first provide an overview of the $x$-axis crossing control scheme, highlighting key findings from the literature that are relevant to our work.This is followed by a description of the root-solving-based formulation, solved using differential correction, and the constrained minimization-based formulation.. Then, the UT-based modification, applicable to either formulation, is introduced. Finally, considerations pertaining to the control frequency and the targeting tolerance are discussed.
@@ 4.1.1 Targeting Event
Consider a spacecraft at $t_{0}$ and state estimate $\boldsymbol{\theta}_{0}\in\mathbb{R}^{6}$. The controller must bring a subset of the propagated state at time $t_{f}$, $\boldsymbol{\psi}(t_{f})\in\mathbb{R}^{n}$ where $n\leq 6$, to an $n$-dimensional ellipsoid from a reference target $\bar{\boldsymbol{\psi}}(\bar{t})\in\mathbb{R}^{n}$ at time $\bar{t}$, with radii $\epsilon_{j,\mathrm{tol}}$ for $j=0,\ldots,n-1$. We denote by targeting event the event that defines $\bar{t}$. While the integration of the dynamics from time $t_{0}$ to $t_{f}$ is typically done in $\mathcal{F}_{\mathrm{I}}$, the targeted state $\boldsymbol{\psi}$ and $\bar{\boldsymbol{\psi}}$ may be components in another frame $\mathcal{F}_{\mathrm{B}}$.
In $x$-axis crossing control, $\bar{t}\neq t_{f}$, as the targeting event is not defined based on propagation time. The non-equivalence of $\bar{t}$ and $t_{f}$ is an important distinction of this control scheme compared to more conventional tracking controllers. The $x$-axis crossing controller gets its name from the fact that the targeting event is defined based on the time where the trajectory crosses the $xz$-plane near the Moon in the Earth-Moon rotating frame, $\mathcal{F}_{\rm EM}$. While this does not exactly coincide with perilune, preliminary experiments have shown that the two events result in no apparent difference in the SK performance, and thus may be used interchangeably; in this work, the targeting event is defined by the perilune.
The optimal impulsive control maneuver $\boldsymbol{u}\in\mathbb{R}^{3}$ that brings the spacecraft to the $n$-dimensional ellipsoid centered at $\bar{\boldsymbol{\psi}}(\bar{t})$ is computed by solving the nonlinear program (NLP)
@@
such that $\left|\boldsymbol{\psi}^{\mathrm{B}}(t_{f})-\bar{\boldsymbol{\psi}}^{\mathrm{B}}(\bar{t})\right|\leq v_{x,\mathrm{tol}},$ (31b)
where the absolute value in constraint (31b) is applied component-wise. The final targeted state $\boldsymbol{\psi}^{\mathrm{B}}(t_{f})$ is the subset of the propagated state at $t_{f}$, $\boldsymbol{\theta}^{\mathrm{B}}(t_{f})$, transformed from the inertial frame to frame $\mathcal{F}_{\mathrm{B}}$, obtained by propagating the nonlinear dynamics with control $\boldsymbol{u}$,
$\boldsymbol{\theta}^{\mathrm{B}}(t_{f})=T^{\mathrm{Inr}}_{\mathrm{B}}\left(\boldsymbol{\theta}_{0}+\begin{bmatrix}\boldsymbol{0}_{3\times 3}\\ \boldsymbol{I}_{3}\end{bmatrix}\boldsymbol{u}+\int_{t_{0}}^{t_{f}}\boldsymbol{f}[\tau,\boldsymbol{\theta}]\mathrm{d}\tau\right),$ (32)
@@
Instead of targeting the immediate perilune, it is common to choose the $N^{\mathrm{th}}$ perilune downstream with $N>1$ to improve the $\Delta V$ performance, subject to the reliability of the predicted state affected by estimation, dynamics, and control execution errors. In this work, $N=7$ is adopted in accordance to previous results on the NRHO [17, 19].
Finally, the $n$ state components to target, the frame in which these state components are represented, and the tolerance on the tightness of the targeting, must be chosen. Previous works have successfully demonstrated that targeting the $x$-component velocity in $\mathcal{F}_{\mathrm{EM}}$, denoted hereafter by $v_{x}^{\mathrm{EM}}$, is sufficient for tracking a baseline NRHO trajectory under reasonable levels of errors and uncertainties [16, 17, 19]. Following their definition, $\mathcal{F}_{\mathrm{B}}$ is $\mathcal{F}_{\mathrm{EM}}$, and $\boldsymbol{\psi}^{\mathrm{B}}=v_{x}^{\mathrm{EM}}$. Then, constraint (31b) is actually
$\left|v_{x}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}(\bar{t})\right|\leq v_{x,\mathrm{tol}},$ (33)
where $v_{x}^{\mathrm{EM}}(t_{f})$ is part of the final state $\boldsymbol{\theta}^{\mathrm{EM}}(t_{f})$ obtained by transforming the solution to the initial value problem (32), and $\bar{v}_{x}^{\mathrm{EM}}(\bar{t})$ is reference $x$-axis velocity in the instantaneous $\mathcal{F}_{\mathrm{EM}}$ at the targeted perilune. One may interpret the $x$-axis crossing controller using $v_{x}^{\mathrm{EM}}$ alone as tracking only the state components necessary for stabilizability. The effect of the tolerance on $v_{x}^{\mathrm{EM}}$ has been investigated by Guzetti et al. [17] using deep space network-based orbit determination, but it is revisited in this work to highlight the controller’s interplay with horizon-based OPNAV.
@@ 4.1.2 Control Location Selection
The control location along the NRHO is parameterized in terms of the instantaneous true anomaly, as shown in Figure 4. The particularly high sensitivity of the dynamics near perilune on the NRHO makes this region sensitive to errors, and is usually considered unsuitable for station-keeping [17]; thus, the control maneuver is to be located at or near the apolune, at a true anomaly of around $180^{\circ}$. Guzetti et al. [17] previously provided results demonstrating a relatively low sensitivity of the maneuver placement at true anomalies between about $160^{\circ}$ and $200^{\circ}$ on the cumulative yearly station-keeping cost. For example, the Gateway’s station-keeping strategy incorporates a single maneuver at $200^{\circ}$ [19], driven by the aforementioned reasons as well as operational considerations. We define the control action true anomaly as the true anomaly about which the controller evaluates the need for control, and if needed, computes the required maneuver, to study its impact on the SK cost for a navigation based on OPNAV.
Figure 4: NRHO shown with instantaneous true anomaly for 10 revolutions, seen from the near-side of the Moon (Earth-side observer) in the Earth-Moon rotating frame. In this view, the spacecraft tracks the orbit anti-clockwise.
@@ 4.2 $x$-axis Crossing Control via Differential Correction
A common solution approach for solving (31) adopted in the literature [16, 17, 19, 36] is differential correction (DC), which involves recasting (31) as an under-determined root-finding problem with a shooting function $F:\mathbb{R}^{3}\to\mathbb{R}$,
$F(\boldsymbol{u})=v_{x}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}(\bar{t}),$ (34)
@@
$t_{0}$, $\boldsymbol{\theta}_{0}$, $t_{f}$, $\bar{v}_{x}^{\mathrm{EM}}$, $v_{x,\mathrm{tol}}$, $\mathrm{maxiter}$

### v2

@@ 3.3.3 Measurement Update
$\boldsymbol{y}_{j}=\boldsymbol{E}\boldsymbol{x}_{j}+\boldsymbol{w}_{j},\;\;\boldsymbol{w}_{j}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{R}_{j}),\;\;\boldsymbol{E}=\begin{bmatrix}\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}\end{bmatrix}$. (34)
@@
$\hat{\boldsymbol{x}}_{j\mid j}$ $=\hat{\boldsymbol{x}}_{j\mid j-1}+\boldsymbol{K}_{j}\left(\boldsymbol{y}_{j}-\boldsymbol{E}\hat{\boldsymbol{x}}_{j\mid j-1}\right),$ (35b)
@@ 3.4 Measurement Collection for Horizon-Based Optical Navigation on Near-Rectilinear Halo Orbit
The choice of the onboard camera specifications directly impacts the number of limb points $m$ detected within an image, and therefore the quality of the position measurement itself. Furthermore, for a given position in space, if the field of view (FOV) is too narrow and the limb of the Moon does not fit within the frame of the image, the measurement may be unusable. The choice of camera FOV is particularly important on the NRHO due to the large variety in distance from the Moon at perilune and apolune. Figure 4 shows the range and the apparent diameter of the Moon as seen along the NRHO across 10 superimposed revolutions.
Figure 4: Range and apparent Moon diameter against osculating true anomaly along NRHO
@@ 3.4.1 Effect of Field of View on Measurement
The least-squares solution of (9) is affected by the number of detected limb points $m$. One way to increase $m$ may be to use a camera with a larger number of pixels on its sensor; however, this requires larger onboard processing overhead. Another tuning knob is the camera’s FOV. A camera with an FOV that exactly matches the apparent diameter of the Moon will contain the largest possible $m$, assuming that the Moon is exactly centered along the boresight. However, in practice, the camera is pointed to the Moon using the state estimate $\hat{\boldsymbol{x}}$. Other operational considerations on attitude may further lead to scenarios where it is desirable to image the Moon without slewing the spacecraft too much.
@@ 3.4.2 Choice of Camera for Horizon-Based OPNAV on NRHO
When choosing the FOV, it is important to keep in mind that for the purpose of autonomous navigation, we seek to have reliable filter estimates near apolune, where station-keeping control maneuvers, when necessary, are executed. On one hand, there is an inherent degradation of observability as the range increases. The worsened observability is manifested in the slow variation of the Moon apparent diameter around apolune, as shown in Figure 4; as the image is taken further away from the Moon, a difference in range $\Delta r$ results in a rapidly diminishing difference in the apparent diameter. The degradation of the measurement quality is reported in previous works, including Figure 9 in [ChristianRobinson2016]. Meanwhile, imaging at perilune, where we have the best observability, also has its downsides: first, from an operational perspective, this may lead to pointing requirements over short periods of time to ensure the Moon is within the frame of the camera. Second, a dead-reckoning prediction from perilune until the control location near apolune may be of insufficient accuracy. Third, even if a measurement collection opportunity may arise around apolune prior to control, a camera with a sufficiently large FOV to capture the full Moon near perilune will only lead to poor measurements with a small number of limb points $m$ near apolune. Thus, we choose an FOV sized to obtain measurements outside of regions along the orbit near perilune; to account for attitude uncertainties, we ensure the FOV is at least twice as large as the Moon’s apparent diameter from where the measurement is acquired.
Note that even with a narrow FOV, the camera may still be used near perilune to provide measurements via alternative, terrain-relative OPNAV methods, such as feature-based [McCabe2020-ck], crater-based navigation [Christian2021-cz] or odometry [Christian2021-qi, De-Vries2022-rt]. In this paper, we choose to focus on horizon-based measurements alone and thereby test its limits in navigating the spacecraft along the NRHO, while fusion with estimates from terrain-relative navigation is possibly left for future studies. For considerations on sensor fusion for cislunar navigation, see e.g. [Vila2025-tl].
@@ 4 Station-Keeping Control
There exists a multitude of controllers proposed for SK on LPOs [Shirobokov2017]. In the context of L1 and L2 halo orbits, a popular SK controller is the so-called $x$-axis crossing control, which has been shown to perform well both in numerical simulations [Davis2017, Guzzetti2017] and in-flight [Cheetham2021]. A particularity about the $x$-axis crossing control scheme is that it has often been formulated as a multivariable root-finding problem, and rarely as a constrained minimization problem [Elango2022]. In this work, we introduce both the root-finding-based and the minimization problem-based formulations, point out their distinctions, and compare their numerical performances. Then, we propose a UT-based modification of the $x$-axis control that utilizes the filter’s covariance estimate together with the state estimate to improve the prediction accuracy of the targeted state, and thereby reduce the cumulative $\Delta V$.
Recent advancements in station-keeping on the NRHO also focused on tackling the phase-deviation issue associated with $x$-axis crossing control [Davis2022, Shimane2024-ep]. To avoid the issue altogether, a model predictive control (MPC) scheme has also been proposed to provide full-state tracking under the same operational assumptions as $x$-axis crossing control [Shimane2025], yielding similar cumulative $\Delta V$ and improved tracking performance. In this work, the conducted numerical experiments do not exceed $\sim\!400$  $\text{\,}\mathrm{d}\mathrm{a}\mathrm{y}\mathrm{s}$, and thus are minimally impacted by phase deviation. The UT-based modification proposed here is applicable to any control scheme, including $x$-axis crossing control schemes with phase-deviation remedies [Davis2022, Shimane2024-ep] or the MPC in [Shimane2025].
In the rest of this Section, we first provide an overview of the $x$-axis crossing control scheme, highlighting key findings from the literature that are relevant to our work. This is followed by a description of the root-solving-based formulation, solved using differential correction, and the constrained minimization-based formulation. Then, the UT-based modification, applicable to either formulation, is introduced. Finally, considerations pertaining to the control frequency and the targeting tolerance are discussed.
@@ 4.1.1 Targeting Event
Consider a spacecraft at $t_{0}$ and state estimate $\boldsymbol{x}_{0}\in\mathbb{R}^{6}$. The controller must bring a subset of the propagated state at time $t_{f}$, $\boldsymbol{\psi}(t_{f})\in\mathbb{R}^{n}$ where $n\leq 6$, to an $n$-dimensional ellipsoid from a reference target $\bar{\boldsymbol{\psi}}(\bar{t}_{f})\in\mathbb{R}^{n}$ at time $\bar{t}_{f}$, with radii $\epsilon_{j,\mathrm{tol}}$ for $j=0,\ldots,n-1$. We denote by targeting event the event that defines $\bar{t}_{f}$. While the integration of the dynamics from time $t_{0}$ to $t_{f}$ is typically done in $\mathcal{F}_{\mathrm{I}}$, the targeted state $\boldsymbol{\psi}$ and $\bar{\boldsymbol{\psi}}$ may be components in another frame $\mathcal{F}_{\mathrm{B}}$.
In $x$-axis crossing control, $\bar{t}_{f}\neq t_{f}$, as the targeting event is not defined based on propagation time. The non-equivalence of $\bar{t}_{f}$ and $t_{f}$ is an important distinction of this control scheme compared to more conventional tracking controllers. The $x$-axis crossing controller gets its name from the fact that the targeting event is defined based on the time where the trajectory crosses the $xz$-plane near the Moon in the Earth-Moon rotating frame, $\mathcal{F}_{\rm EM}$. While this does not exactly coincide with perilune, preliminary experiments have shown that the two events result in no apparent difference in the SK performance, and thus may be used interchangeably; in this work, the targeting event is defined by the perilune.
The optimal impulsive control maneuver $\boldsymbol{u}\in\mathbb{R}^{3}$ that brings the spacecraft to the $n$-dimensional ellipsoid centered at $\bar{\boldsymbol{\psi}}(\bar{t}_{f})$ is computed by solving the nonlinear program (NLP)
@@
such that $\left|\boldsymbol{\psi}^{\mathrm{B}}(t_{f})-\bar{\boldsymbol{\psi}}^{\mathrm{B}}(\bar{t}_{f})\right|\leq v_{x,\mathrm{tol}},$ (36b)
where the absolute value in constraint (36b) is applied component-wise. The final targeted state $\boldsymbol{\psi}^{\mathrm{B}}(t_{f})$ is the subset of the propagated state at $t_{f}$, $\boldsymbol{x}^{\mathrm{B}}(t_{f})$, transformed from the inertial frame to frame $\mathcal{F}_{\mathrm{B}}$, obtained by propagating the nonlinear dynamics with control $\boldsymbol{u}$,
$\boldsymbol{x}^{\mathrm{B}}(t_{f})=T^{\mathrm{Inr}}_{\mathrm{B}}\left(\boldsymbol{x}_{0}+\begin{bmatrix}\boldsymbol{0}_{3\times 3}\\ \boldsymbol{I}_{3}\end{bmatrix}\boldsymbol{u}+\int_{t_{0}}^{t_{f}}\boldsymbol{f}[\tau,\boldsymbol{x}]\mathrm{d}\tau\right),$ (37)
@@
Instead of targeting the immediate perilune, it is common to choose the $N^{\mathrm{th}}$ perilune downstream with $N>1$ to improve the $\Delta V$ performance, subject to the reliability of the predicted state affected by estimation, dynamics, and control execution errors. In this work, $N=7$ is adopted in accordance to previous results on the NRHO [Guzzetti2017, Davis2022].
Finally, the $n$ state components to target, the frame in which these state components are represented, and the tolerance on the tightness of the targeting, must be chosen. Previous works have successfully demonstrated that targeting the $x$-component velocity in $\mathcal{F}_{\mathrm{EM}}$, denoted hereafter by $v_{x}^{\mathrm{EM}}$, is sufficient for tracking a baseline NRHO trajectory under reasonable levels of errors and uncertainties [Davis2017, Guzzetti2017, Davis2022]. Following their definition, $\mathcal{F}_{\mathrm{B}}$ is $\mathcal{F}_{\mathrm{EM}}$, and $\boldsymbol{\psi}^{\mathrm{B}}=v_{x}^{\mathrm{EM}}$. Then, constraint (36b) is actually
$\left|v_{x}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}(\bar{t}_{f})\right|\leq v_{x,\mathrm{tol}},$ (38)
where $v_{x}^{\mathrm{EM}}(t_{f})$ is part of the final state $\boldsymbol{x}^{\mathrm{EM}}(t_{f})$ obtained by transforming the solution to the initial value problem (37), and $\bar{v}_{x}^{\mathrm{EM}}(\bar{t}_{f})$ is reference $x$-axis velocity in the instantaneous $\mathcal{F}_{\mathrm{EM}}$ at the targeted perilune. One may interpret the $x$-axis crossing controller using $v_{x}^{\mathrm{EM}}$ alone as tracking only the state components necessary for stabilizability. The effect of the tolerance on $v_{x}^{\mathrm{EM}}$ has been investigated by Guzetti et al. [Guzzetti2017] using deep space network-based orbit determination, but it is revisited in this work to highlight the controller’s interplay with horizon-based OPNAV.
@@ 4.1.2 Control Location Selection
The control location along the NRHO is parameterized in terms of the instantaneous true anomaly $\theta$, calculated assuming a Moon-centered restricted two-body problem,
$\theta=\operatorname{atan2}{\left(hv_{r},h^{2}/r-\mu\right)},\,\,\,h=\|\boldsymbol{r}\times\boldsymbol{v}\|_{2},\,\,\,v_{r}=\dfrac{\boldsymbol{r}\cdot\boldsymbol{v}}{r},$ (39)
as shown in Figure 5. In the Figure, the black arrows indicate the direciton of motion. The particularly high sensitivity of the dynamics near perilune on the NRHO makes this region sensitive to errors, and is usually considered unsuitable for station-keeping [Guzzetti2017]; thus, the control maneuver is to be located at or near the apolune, at a true anomaly of around $180^{\circ}$. Guzetti et al. [Guzzetti2017] previously provided results demonstrating a relatively low sensitivity of the maneuver placement at true anomalies between about $160^{\circ}$ and $200^{\circ}$ on the cumulative yearly station-keeping cost. For example, the Gateway’s station-keeping strategy incorporates a single maneuver at $200^{\circ}$ [Davis2022], driven by the aforementioned reasons as well as operational considerations. We define the control action true anomaly as the true anomaly about which the controller evaluates the need for control, and if needed, computes the required maneuver, to study its impact on the SK cost for a navigation based on OPNAV.
Figure 5: NRHO over 10 revolutions in Moon-centered Earth-Moon rotating frame
@@ 4.2 $x$-axis Crossing Control via Differential Correction
A common solution approach for solving (36) adopted in the literature [Davis2017, Guzzetti2017, Davis2022, Muralidharan2022] is differential correction (DC), which involves recasting (36) as an under-determined root-finding problem with a shooting function $F:\mathbb{R}^{3}\to\mathbb{R}$,
$F(\boldsymbol{u})=v_{x}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}(\bar{t}_{f}),$ (40)
@@
$t_{0}$, $\boldsymbol{x}_{0}$, $t_{f}$, $\bar{v}_{x}^{\mathrm{EM}}$, $v_{x,\mathrm{tol}}$, $\mathrm{maxiter}$